OpenStudy (anonymous):

For the reaction S + O2 yields SO2 change in heat is -296 kj per mole of SO2 formed. Calculate the quantity of heat released when 1.00 g of sulfur is burned in oxygen

OpenStudy (aaronq):

\(\sf \large heat~released=\Delta H^o*moles\) and \(\sf moles=\dfrac{mass}{molar~mass}\) therefore: \(\sf \large heat~released=\Delta H^o*\dfrac{mass}{molar~mass}\)